OpenStudy (anonymous):

In set-builder notation, how do you write the solutions of the inequality? 3x + 10 > with line under it 4 x - 2 x 2 x - 2 x 2

OpenStudy (anonymous):

\[3x+10\geq 4\] \[3x\geq -6\] \[x\geq -2\]